OpenStudy (anonymous):

solve my substitution (medal and fan) -x - y - z = -8 -4x + 4y + 5z = 7 2x + 2z = 4

OpenStudy (amoodarya):

rewrite it -x - y - z = -8 -4x + 4y + 5z = 7 x + z = 2 from 3rd x=2-z so put in 1st and 2nd one

OpenStudy (anonymous):

im still confused :(

OpenStudy (amoodarya):

i divided the 3rd equation by 2 so have x+y=4/2=2

OpenStudy (amoodarya):

let start from the begining -x - y - z = -8 -4x + 4y + 5z = 7 x + z = 2 from 1st equation -x-y-z=-8 -x=-8+y+z so multiply by -1 x=8-y-z

OpenStudy (amoodarya):

put x=8-y-z in other equations -4x + 4y + 5z = 7 x + z = 2

OpenStudy (amoodarya):

-4(8-y-z) + 4y + 5z = 7 (8-y-z)+ z = 2

OpenStudy (amoodarya):

can you go on ?

OpenStudy (amoodarya):

-4(8-y-z) + 4y + 5z = 7 (8-y-z)+ z = 2 --->8-y=2 --->y=6 put that in above

OpenStudy (anonymous):

yeah! thankss